If the electronegativity difference between two atoms A and B is 2.0, then the percentage distribution character in the molecule is


(A)        54%


(B)        46%


(C)        23%


(D)        72%


 


Grade:11

1 Answers

Shane Macguire
10 years ago

The percentage ionic character is related to electronegativity difference between the two atoms. It is given by Hannay and Smith relationship.

 

        %age ionic character.

 

                = 16 (XA − XB) + 3.5 (XA − XB)2

 

        Here, XA and XB are elecrtronegativities of two atoms A and B

 

        \ % ionic character = 16(2) + 3.5(2)2 = 46%

 

        or covalent character = 100 – 46 = 54%